SPECTROSCOPIC ANALYSIS OF THE A AND 3 ^{1}Σ^{+} STATES OF ^{39}K^{85}Rb

JIN-TAE KIM, Department of Photonic Engineering, Chosun University, Gwangju, 501-759, Korea; YONGHOON LEE, Department of Chemistry, Mokpo National University, Jeonnam 534-729, Korea; BONGSOO KIM, Department of Chemistry, KAIST, Daejeon, 305-701, Korea; DAJUN WANG, Department of Physics,The Chinese University of Hong Kong, Shatin, Hong Kong; PHILLIP L. GOULD, \underline{EDWARD E. EYLER}, and WILLIAM C. STWALLEY, Department of Physics, University of Connecticut, Storrs, CT 06269-3046, USA.

The vibrational levels of the mutually strongly perturbed 1 ^{1}Π − 2 ^{3}Σ^{+} − b ^{3}Π states of ^{39}K^{85}Rb^{a} were previously assigned by using a combination of molecular beam and ultracold molecule (UM) excitation spectra, the latter generated from levels of the a ^{3}Σ^{+} state formed following photoassociation to a level of the 3(0^{−}) state. In a follow-on study using the same techniques, we have now successfully assigned high vibrational levels of the A and 3 ^{1}Σ^{+} states from the excitation spectra of UMs formed by PA to both 3(0^{+}) and 3(0^{−}) levels. The ^{1}Σ^{+} states are absent in the UM spectra for levels formed by the 3(0^{−}) PA level. This absence has been explained by considering Hund’s case (c) selection rules and the transition dipole moment calculations by Kotochigova et al.^{b} between the upper excited A ^{1}Σ^{+}(2(0^{+})) state and the three Ω components at the ground state dissociation limit. This work is supported by NRF in Korea (Grant 2009-0085319 at Chosun and Grants 2011-0001335 and 2011-0020419 at KAIST) and by NSF (Grant PHY-0855613) and AFOSR (Grant MURI FA9550-09-1-588) in the US.

aJ. T. Kim, Y. Lee, B. Kim, D. Wang, W. C. Stwalley, P. L. Gould, and E. E. Eyler, Phys. Chem. Chem. Phys. 13, 18755 (2011).

bS. Kotochigova, E. Tiesinga, and P. S. Julienne, Phys. Rev. A 68, 022501 (2003).